# SpecDec++: Boosting Speculative Decoding via Adaptive Candidate Lengths

### v2

@@ Abstract
Speculative decoding reduces the inference latency of a target large language model via utilizing a smaller and faster draft model. Its performance depends on a hyperparameter $K$ — the candidate length, i.e., the number of candidate tokens for the target model to verify in each round. However, previous methods often use simple heuristics to choose $K$, which may result in sub-optimal performance. We study the choice of the candidate length $K$ and formulate it as a Markov Decision Process. We theoretically show that the optimal policy of this Markov decision process takes the form of a threshold policy, i.e., the current speculation should stop and be verified when the probability of getting a rejection exceeds a threshold value. Motivated by this theory, we propose SpecDec++, an enhanced version of speculative decoding that adaptively determines the candidate length on the fly. We augment the draft model with a trained acceptance prediction head to predict the conditional acceptance probability of the candidate tokens. SpecDec++ will stop the current speculation when the predicted probability that at least one token gets rejected exceeds a threshold. We implement SpecDec++ and apply it to the llama-2-chat 7B & 70B model pair. Our adaptive method achieves a 2.04x speedup on the Alpaca dataset (an additional 7.2% improvement over the baseline speculative decoding). On the GSM8K and HumanEval datasets, our method achieves a 2.26x speedup (9.4% improvement) and 2.23x speedup (11.1% improvement), respectively.
@@ 1 Introduction
Current state-of-the-art Large Language Models (LLMs) have demonstrated extraordinary capabilities in various language tasks and have shown early signs of artificial general intelligence [1, 3, 40, 41, 42]. As the top-performing LLMs often have more than a hundred billion parameters, there is an increasing demand for serving such huge models efficiently. To decrease the inference latency, motivated by speculative execution techniques in processors, speculative decoding [8, 24] incorporates a draft model, which is smaller and faster, as the speculator for the target model, which is the large language model we want to accelerate. Given the current prefix, the draft model first auto-regressively generates $K$ tokens, taking substantially less time than it would take the target model. The target model computes their log probabilities in parallel and then sequentially determines whether each token is accepted or not. Following the first rejected token (if any), the algorithm discards the remaining tokens and corrects the rejected token with a fresh sample from a modified distribution. If all tokens are accepted, a new token is sampled from the next-token probability given by the target model and appended to the sequence of accepted tokens, and then the process moves forward. Such draft-verify-correct loops continue until the desired output is fully generated.
@@
Leviathan et al. [24] studied the problem of choosing the hyperparameter $K$ under the assumption that the acceptance rates of all the candidate tokens are constant. The authors showed that there exists one constant $K$ that can maximize the speedup. However, such an assumption is unrealistic and does not approximate real-world cases well. Whether the draft model and the target model align well depends on the hardness of predicting the next token. Intuitively, when the next token is unambiguous from the prefix, the draft model and the target model align well, which means the acceptance probability of the current candidate token is large compared to other cases.
@@
In this work, we aim to boost the performance of speculative decoding by adaptively choosing the candidate length $K$ for each round. We formalize the adaptive decision-making of $K$ for speculative decoding as a Markov Decision Process (MDP). The decision to make at each timestep is whether or not to stop the current speculation round and submit the candidate tokens to the target model for verification and correction. The objective is to minimize the total inference time taken to generate a full response. Theoretically, we show that the optimal policy takes the form of a threshold policy, i.e., it is optimal to stop the speculation round whenever the probability of existing at least one rejected token in the candidates exceeds a threshold.
Inspired by the theory, we propose SpecDec++, an enhanced version of speculative decoding that adaptively determines the candidate length on the fly. First, we train an acceptance prediction head on top of the draft model to predict the acceptance probability of the candidate token. Training such an acceptance prediction head has two challenges: (1) there will be a severe class imbalance problem, e.g., most tokens generated by the draft model will have a high probability of acceptance, depending on how well the two models align; (2) the input sequence to the model contains mostly tokens from the target model and only a fraction of tokens generated by the draft model, so the training signal is sparse. To overcome the two challenges, we adopt a weighted Binary Cross-Entropy loss to address the class imbalance problem, and we adapt the random masking idea from BERT [14] to randomly mix tokens from the target model and the draft model to increase training efficiency.
At inference time, we opt to stop the current speculation round when the predicted probability of the existence of a rejected token exceeds a constant stopping threshold. The procedure is illustrated in Figure 2. We implement SpecDec++ and apply it to llama-2-chat 7B & 70B model pair. Our adaptive method achieves a 2.04x speedup compared with the 1.90x speedup of the baseline speculative decoding method on the Alpaca dataset (an additional 7.2% improvement). On the easier GSM8K and HumanEval datasets, our method improves the baseline from 2.07x to 2.26x speedup (9.4% improvement) and from 2.00x to 2.23x speedup (11.1% improvement), respectively.
@@
  • We formalize the dynamic choice of candidate length in speculative decoding as a Markov Decision Process (MDP). We theoretically show that when the probability that at least one token gets rejected exceeds a threshold, the optimal action is to stop the speculation and submit it for verification.
@@
  • We propose SpecDec++, an enhanced version of speculative decoding that adaptively determines the candidate length on the fly. We adopt a weighted loss and a token-mixing method to efficiently train the prediction head and use it for dynamic decision-making in the decoding process.
@@
  • Empirically, our method achieves an additional 7.2%, 9.4%, and 11.1% improvement over the baseline speculative decoding on the Alpaca, HumanEval, and GSM8K datasets, respectively.
2 Background
Rejection Sampling. If we want to sample from a target discrete distribution $p(x)$, we first sample $x$ from a draft distribution $q(x)$. We accept the sample $x$ with probability $\min(1,\frac{p(x)}{q(x)})$; otherwise we replace it with a sample from the modified distribution $\mathrm{Norm}[(p-q)_{+}]$, where $z_{+}=\max(z,0)$ is the positive part of $z$ and $\mathrm{Norm}[f]=\frac{f(\cdot)}{\sum_{x}f(x)}$ normalizes a function $f$ to make it a proper probability distribution. The proof of the unbiasedness of rejection sampling can be found in [8].
@@ 2 Background
Speculative Decoding. Speculative decoding extends to the auto-regressive generation scenarios by chaining $K$ rejection sampling procedures together. To auto-regressively generate a sequence from $p(\cdot\mid x_{\text{prefix}})$, we first generates $K$ candidate tokens $(y_{1},y_{2},\dots,y_{K})$ from $q(\cdot\mid x_{\text{prefix}})$
$y_{i}\sim q(Y_{i}\mid x_{\text{prefix}},y_{1},\dots,y_{i-1}),\quad i=1,2,\dots,K.$
@@
Next, we sequentially check if each $y_{i}$ is accepted or not. If there is any rejection, we replace the first rejected token with a fresh sample from the corresponding modified probability distribution and discard the subsequent tokens.
@@
The key practical consideration is that the probabilities of the candidate tokens $p(y_{i}\mid x_{\text{prefix}},y_{1},\dots,y_{i-1})$ can be calculated in parallel by the target model with no additional overhead, as the forward time is bottlenecked by the memory operations [32]. For completeness, the speculative decoding algorithm is stated in Algorithm 1.
Algorithm 1 Speculative Decoding [8, 24]
  for $i=1$ to $K$  do
     Compute $q_{i}=q(\cdot\mid x_{\text{prefix}},y_{1},\dots,y_{i-1})$.
     Sample $y_{i}\sim q_{i}$.
  end for
   Sample $r_{1},\dots,r_{K}$ with $r_{i}\sim\mathrm{Unif}[0,1]$, $i=1,\dots,K$.
  if $n<K$ then
  else
     Sample $y^{\prime}$ from $p_{K+1}$
  end if
  Return $x_{\text{prefix}},y_{1},\dots,y_{n},y^{\prime}$
Inference Time of Speculative Decoding.
Our objective is to minimize the total inference time, which consists of
@@
where $t_{\text{draft}}$ and $t_{\text{target}}$ are the time needed for one forward pass and $N_{\text{draft}}$ and $N_{\text{target}}$ are the total number of forward passes of the draft model and the target model, respectively. Equation (2.1) holds under the implicit assumption that the forward passes of each of the models take constant time, which is true when we have enough computational resources to support the increased concurrency when the length of the input sequence grows [24]. We empirically verify that Equation (2.1) holds in our setting; see Section 4.2.
Let $N$ be the number of the final generated tokens. $N$ is a random variable inherent to the target model and the initial prompt, independent of the draft model and the number of candidate tokens $K$ of each round we choose. Let $N_{\text{discarded}}$ be the number of total discarded tokens. Then we have the following identity for Algorithm 1
$N_{\text{draft}}+N_{\text{target}}=N+N_{\text{discarded}}.$
Therefore, $T_{\text{total}}$ can be written as
@@
To minimize the total inference time, we are required to trade-off between two objectives: minimizing the number of the discarded tokens $N_{\text{discarded}}$ and minimizing the number of forward passes of the target model $N_{\text{target}}$. The two objectives conflict with each other, as a larger $K$ will incur more discarded tokens but less number of forward passes of the target model. Equation (2.2) states that the total cost is the weighted sum of the two and the weights are given by $t_{\text{draft}}$ and $(t_{\text{target}}-t_{\text{draft}})$.
We focus on two metrics: (1) discard rate $N_{\text{discarded}}/N$, which measures the average number of discarded tokens per one generated token, and (1) verification rate $N_{\text{target}}/N$, which measures the average number of the forward calls of the target model per one generated token.
2.1 A Motivating Example: Oracle Performances of Greedy Speculative Decoding
@@
To empirically verify this, we perform a preliminary experiment with the same setting in Section 4, where we use all the prompts in the Alpaca dataset and calculate the set of disagreement $S$ for each prompt with the llama-2-chat-7B/llama-2-chat-70B model pair. The results show that the average $N_{\text{target}}/N=0.164\pm 0.078$ and the corresponding oracle throughput is $27.06\pm 4.13$ tokens/second (2.92x speedup) according to Equation (2.3) with the empirical value of $(t_{\text{target}},t_{\text{draft}})$ reported in Section 4.2. In comparison, the average throughput for the target model without speculative decoding is 9.26 tokens/second, while speculative decoding with the best fixed $K$ gives 17.58 tokens/second (1.90x speedup) (Section 4). We see that there is a huge potential in adaptively tuning the candidate lengths.
3 SpecDec++: Theory and Algorithm
@@ 3.1 Speculative Decoding as Markov Decision Processes
Actions. Given the current state $(x_{\text{prefix}},(Y_{1},\dots,Y_{k}))$, the decision to make is whether or not to end the current speculation round and submit the candidate tokens to the target model for verification. We denote the current action by $a\in\{{\mathrm{stop}},{\mathrm{continue}}\}$ as the choice of stopping or continuing the current speculation round.  [^1] In practice, when $Y_{k+1}$ is EOS (the special token denoting the end of sequence) or when the total length hits the maximal generation length, we manually set $a={\mathrm{stop}}$.
We note that in an extended MDP setting, we can include the draft probability $q_{k+1}$ for the token $Y_{k+1}$ as a part of the current action. Finetuning the draft model to align better with the target model can be viewed as an offline policy optimization algorithm that will likely improve the performance. And it has been studied in previous work, e.g. DistillSpec [54] and Medusa [7]. In the paper, we consider the draft probability $q_{k+1}$ as given by the draft model and do not optimize $q_{k+1}$.
@@
  • When $a={\mathrm{continue}}$, the next state $s^{\prime}$ is simply $(x_{\text{prefix}},(Y_{1},\dots,Y_{k},Y_{k+1}))$.
  • When $a={\mathrm{stop}}$, the candidate tokens $(Y_{1},\dots,Y_{k+1})$ are verified via speculative decoding (Algorithm 1). Let $n$ be the number of the accepted tokens. Let $y^{\prime}$ be the replaced token when $n<k+1$ or the fresh token from the next-token distribution given by the target model when $n=k+1$. The next state $s^{\prime}=(x_{\text{prefix}}^{\prime},\varnothing)$ with the new prefix $x_{\text{prefix}}^{\prime}=(x_{\text{prefix}},y_{1},\dots,y_{n},y^{\prime})$ being the concatenation of the previous prefix and the newly generated tokens.
Immediate Costs. According to Equation (2.2), let $c_{1}=t_{\text{draft}}$ and $c_{2}=(t_{\text{target}}-t_{\text{draft}})$. We can define the immediate cost as the following
$c(s,{\mathrm{continue}},s^{\prime})=\mathbb{I}(\exists 1\leq i\leq k+1,Y_{i}\text{ is rejected})\cdot c_{1},$
$c(s,{\mathrm{stop}},s^{\prime})=\mathbb{I}(\exists 1\leq i\leq k+1,Y_{i}\text{is rejected})\cdot c_{1}+c_{2}.$
For both cases, we suffer a loss $c_{1}$ if the current candidate token $Y_{k+1}$ is discarded, which happens if there exists any candidate token $Y_{i}$ $(1\leq i\leq k+1)$ that is rejected. If we choose to stop at the current step, we suffer an additional cost $c_{2}$ corresponding to the extra inference time of the target model.
Note that different from the traditional MDP setting when the reward/cost is immediately available to the learner, our setting is more related to the delayed feedback setting [18, 23, 50, 10], where the cost is only available after the candidate tokens are submitted to the target model for verification.
@@
the expected total cost of ${\mathrm{stop}}$ is smaller than the expected total cost of ${\mathrm{continue}}$, where $\Delta=\Delta(\pi,x_{\text{prompt}},p,q,c_{1},c_{2})$ is a problem-specific constant.
We defer the proof of Theorem 3.1 to Appendix D.
3.2 SpecDec++
Theorem 3.1 provides a sufficient condition for us to stop the current round of speculation and call the target model to verify the candidate tokens. Motivated by Theorem 3.1, we propose SpecDec++, an adaptive speculative decoding algorithm that utilizes an additional prediction head to determine whether or not to stop the current speculation round.
SpecDec++ incorporates an additional prediction head $f_{\theta}$ on top of the draft model that predicts the conditional probability

### v3

@@ Abstract
Speculative decoding reduces the inference latency of a target large language model via utilizing a smaller and faster draft model. Its performance depends on a hyperparameter $K$ — the candidate length, i.e., the number of candidate tokens for the target model to verify in each round. However, previous methods often use simple heuristics to choose $K$, which may result in sub-optimal performance. We study the choice of the candidate length $K$ and formulate it as a Markov Decision Process. We theoretically show that the optimal policy of this Markov decision process takes the form of a threshold policy, i.e., the current speculation should stop and be verified when the probability of getting a rejection exceeds a threshold value. Motivated by this theory, we propose SpecDec++, an enhanced version of speculative decoding that adaptively determines the candidate length on the fly. We augment the draft model with a trained acceptance prediction head to predict the conditional acceptance probability of the candidate tokens. SpecDec++ will stop the current speculation when the predicted probability that at least one token gets rejected exceeds a threshold. We implement SpecDec++ and apply it to the llama-2-chat 7B & 70B model pair. Our adaptive method achieves a 2.04x speedup on the Alpaca dataset (7.2% improvement over the baseline speculative decoding). On the GSM8K and HumanEval datasets, our method achieves a 2.26x speedup (9.4% improvement) and 2.23x speedup (11.1% improvement), respectively. The code of this paper is available at https://github.com/Kaffaljidhmah2/SpecDec_pp.
@@ 1 Introduction
Current state-of-the-art Large Language Models (LLMs) have demonstrated extraordinary capabilities in various language tasks and have shown early signs of artificial general intelligence (Achiam et al., 2023; Anil et al., 2023; Team et al., 2023; Touvron et al., 2023a; b). As the top-performing LLMs often have hundreds of billions of parameters and extremely long context windows, there is an increasing demand for serving such huge models efficiently.
To decrease the inference latency, motivated by speculative execution techniques in processors, speculative decoding (Chen et al., 2023a; Leviathan et al., 2023) incorporates a draft model, which is smaller and faster, as the speculator for the target model, which is the large language model we want to accelerate. Given the current prefix, the draft model first auto-regressively generates $K$ tokens, taking substantially less time than it would take the target model. The target model computes their log probabilities in parallel and then sequentially determines whether each token is accepted or not. Following the first rejected token (if any), the algorithm discards the remaining tokens and corrects the rejected token with a fresh sample from a modified distribution. If all tokens are accepted, a new token is sampled from the next-token probability given by the target model and appended to the sequence of accepted tokens, and then the process moves forward. Such draft-verify-correct loops continue until the desired output is fully generated.
@@
Leviathan et al. (2023) studied the problem of choosing the hyperparameter $K$ under the assumption that the acceptance rates of all the candidate tokens are constant. The authors showed that there exists one constant $K$ that can maximize the speedup. However, such an assumption is unrealistic and does not approximate real-world cases well. Whether the draft model and the target model align well depends on the hardness of predicting the next token. Intuitively, when the next token is unambiguous from the prefix, the draft model and the target model align well, which means the acceptance probability of the current candidate token is large compared to other cases.
@@
In this work, we aim to boost the performance of any speculative decoding algorithm by adaptively choosing the candidate length $K$ for each round. We first formalize the adaptive decision-making of $K$ for speculative decoding as a Markov Decision Process (MDP). The decision to make at each timestep is whether or not to stop the current speculation round and submit the candidate tokens to the target model for verification and correction. The objective is to minimize the total inference time taken to generate a full response. Theoretically, we show that the optimal policy takes the form of a threshold policy, i.e., it is optimal to stop the speculation round whenever the probability of existing at least one rejected token in the candidates exceeds a threshold.
Inspired by the theory, we propose SpecDec++, an enhanced version of speculative decoding that adaptively determines the candidate length on the fly. First, we train an acceptance prediction head on top of the draft model to predict the acceptance probability of the candidate token. Training such an acceptance prediction head has two challenges: (1) there will be a severe class imbalance problem, e.g., most tokens generated by the draft model will have a high probability of acceptance, depending on how well the two models align; (2) the input sequence to the model contains mostly tokens from the target model and only a fraction of tokens generated by the draft model, so the training signal is sparse. To overcome the two challenges, we adopt a weighted Binary Cross-Entropy loss to address the class imbalance problem, and we develop a token mixing approach by randomly mixing tokens from the target model and the draft model to increase training efficiency.
At inference time, we opt to stop the current speculation round when the predicted probability of the existence of a rejected token exceeds a constant stopping threshold. The procedure is illustrated in Figure 2. To validate the effectiveness of our proposed improvement technique, we choose the simplest implementation of speculative decoding as the baseline and augment it with SpecDec++. When evaluating on llama-2-chat 7B & 70B model pair, our adaptive method achieves a 2.04x speedup compared with the 1.90x speedup of the baseline speculative decoding method on the Alpaca dataset. On the easier GSM8K and HumanEval datasets, our method boosts the baseline from 2.07x to 2.26x speedup and from 2.00x to 2.23x speedup , respectively.
@@
  • We formalize the dynamic choice of candidate length in speculative decoding as a Markov Decision Process (MDP) and conduct a rigorous study on the inference time. We theoretically show that when the probability that at least one token gets rejected exceeds a threshold, the optimal action is to stop the speculation and submit it for verification.
@@
  • We propose SpecDec++, an enhanced version of speculative decoding that adaptively determines the candidate length on the fly. We develop a weighted binary cross-entropy loss and a token mixing method to efficiently train the prediction head and use it for dynamic decision-making in the decoding process.
@@
  • We validate the effectiveness of SpecDec++ with a simple baseline implementation. Our method achieves an additional 7.2%, 9.4%, and 11.1% improvement over the baseline speculative decoding on the Alpaca, HumanEval, and GSM8K datasets, respectively.
2 Related Work
Improvements on Speculative Decoding. Since the proposal of speculative decoding, people have been improving the algorithm from different perspectives, for example, (1) making the draft model align better with the target model (Zhou et al., 2024; Agarwal et al., 2024; Liu et al., 2023), (2) building smaller draft models or merging draft models into the target model (e.g. early-exiting) (Miao et al., 2023; Liu et al., 2024; Yang et al., 2023b; Bae et al., 2023; Zhang et al., 2024; Monea et al., 2023; Chen et al., 2023b), and (3) building a heirachical system of speculative decoding (Spector & Re, 2023; Sun et al., 2024a).
In contrast, our work focuses on the theoretical properties of the candidate length selection, and our improvement is achieved through algorithmic improvements that are independent of system-level and hardware-level configurations and orthogonal to architectural or system-level improvements. This means SpecDec++ can be plugged into any implementation of speculative decoding as long as it adopts a form of draft models, e.g., EAGLE (Li et al., 2024b). Furthermore, it can be readily combined with other system and hardware improvements.
Medusa-like Methods. Several studies improve speculative decoding by abandoning the auto-regressive draft model, including blockwise parallel sampling (Stern et al., 2018) and the popular Medusa (Cai et al., 2024). However, these approaches typically adopt a different method for verifying the candidate tokens, and the generated tokens may deviate from the target model’s distribution under the general stochastic sampling setting. Therefore, we choose not to compare against this line of methods in our paper.
Heuristic Candidate Length Selection Methods. Leviathan et al. (2023) make the i.i.d. assumption on the acceptance probabilities of the candidate tokens and theoretically derive the optimal choice of $K$. Besides, Liu et al. (2024) and Kim et al. (2024) adopt a simple heuristic that ends the speculation if the confidence of the current draft token distribution falls below a threshold. Xu et al. (2023) uses the cumulative product of the confidences and extends to token trees. We include a discussion on why simple heuristics like confidence or entropy may lead to sub-optimal performance in Appendix B.1.
In comparison, our work systematically studies the candidate length selection within the theoretical MDP framework and uses the cumulative product of our trained prediction head to determine the end of the speculation. Due to space limit, please see Appendix B for an extended related work section.
3 Inference Time Analysis of Speculative Decoding
3.1 Background of Speculative Decoding
@@
To auto-regressively generate a sequence from $p(\cdot\mid x_{\text{prefix}})$ using speculative decoding, we first generate $K$ candidate tokens $(y_{1},y_{2},\dots,y_{K})$ from $q(\cdot\mid x_{\text{prefix}})$
$y_{i}\sim q(Y_{i}\mid x_{\text{prefix}},y_{1},\dots,y_{i-1}),\quad i=1,2,\dots,K,$
@@
We refer to $K$ as the candidate length, i.e., the number of candidate tokens for this round. Next, we sequentially check if each $y_{i}$ is accepted or not. If there is any rejection, we replace the first rejected token with a fresh sample from the corresponding modified probability distribution and discard the subsequent tokens. For completeness, the details of the speculative decoding algorithm are stated in Appendix C.
@@
The key practical consideration is that the probabilities of the candidate tokens $p(y_{i}\mid x_{\text{prefix}},y_{1},\dots,y_{i-1})$ can be calculated in parallel by the target model with no additional overhead, as the forward time is bottlenecked by the memory operations (Pope et al., 2023).
3.2 Inference Time Decomposition of Speculative Decoding
Our objective is to minimize the total inference time, which satisfies
@@ 3.2 Inference Time Decomposition of Speculative Decoding
where $t_{\text{draft}}$ and $t_{\text{target}}$ are the time needed for one forward pass and $N_{\text{draft}}$ and $N_{\text{target}}$ are the total number of forward passes of the draft model and the target model, respectively. Equation (3.1) holds under the implicit assumption that the forward passes of each of the models take constant time, which is true when we have enough computational resources to support the increased concurrency when the length of the input sequence grows (Leviathan et al., 2023). We empirically verify that Equation (3.1) holds in our setting; see Section 5.2.
Let $N$ be the number of the final generated tokens. Note that $N$ is a random variable inherent to the target model and the initial prompt, independent of the draft model and the number of candidate tokens $K$ of each round we choose. Let $N_{\text{discarded}}$ be the number of total discarded tokens. By the fact that $N_{\text{draft}}+N_{\text{target}}=N+N_{\text{discarded}}$, we have the following lemma.
Lemma 3.1.
The total inference time of any speculative decoding algorithm $T_{\text{total}}$ can be decomposed as
@@
To minimize the total inference time, we are required to trade-off between two objectives: minimizing the number of the discarded tokens $N_{\text{discarded}}$ and minimizing the number of forward passes of the target model $N_{\text{target}}$. The two objectives conflict with each other, as a larger candidate length $K$ will incur more discarded tokens but less number of forward passes of the target model. Equation (3.2) states that the total cost is the weighted sum of the two and the weights are given by $t_{\text{draft}}$ and $(t_{\text{target}}-t_{\text{draft}})$.
4 SpecDec++: Theory and Algorithm
4.1 A Motivating Example: Oracle Performance of Greedy Speculative Decoding
@@
Empirical implication. We perform a preliminary study where we use all the prompts in the Alpaca dataset and calculate the set of disagreement $S$ for each prompt with the llama-2-chat-7B/llama-2-chat-70B model pair. The results show that the average $N_{\text{target}}/N=0.164\pm 0.078$ and the corresponding oracle throughput is $27.06\pm 4.13$ tokens/second (2.92x speedup) in the setting of Section 5. In comparison, the average throughput for the target model without speculative decoding is 9.26 tokens/second, while speculative decoding with the best fixed $K$ gives 17.58 tokens/second (1.90x speedup) (Section 5). We can see a huge potential in adaptively tuning the candidate lengths, which motivates our subsequent study on stochastic settings and the development of SpecDec++.
@@ 4.2 Speculative Decoding as Markov Decision Processes
Actions. Given the current state $(x_{\text{prefix}},(Y_{1},\dots,Y_{k}))$, the decision to make is whether or not to end the current speculation round and submit the candidate tokens to the target model for verification. We denote the current action by $a\in\{{\texttt{stop}},{\texttt{continue}}\}$ as the choice of stopping or continuing the current speculation round.  [^1] In practice, when $Y_{k+1}$ is EOS (the special token denoting the end of sequence) or when the total length hits the maximal generation length, we manually set $a={\texttt{stop}}$.
@@
  • When $a={\texttt{continue}}$, the next state $s^{\prime}$ is simply $(x_{\text{prefix}},(Y_{1},\dots,Y_{k},Y_{k+1}))$.
  • When $a={\texttt{stop}}$, the candidate tokens $(Y_{1},\dots,Y_{k+1})$ are verified via speculative decoding (Algorithm 2). Let $n$ be the number of the accepted tokens. Let $y^{\prime}$ be the replaced token when $n<k+1$ or the fresh token from the next-token distribution given by the target model when $n=k+1$. The next state $s^{\prime}=(x_{\text{prefix}}^{\prime},\varnothing)$ with the new prefix $x_{\text{prefix}}^{\prime}=(x_{\text{prefix}},y_{1},\dots,y_{n},y^{\prime})$ being the concatenation of the previous prefix and the newly generated tokens.
Our next theorem provides a simple way to set the immediate cost so that the cumulative cost of the MDP matches the total inference time of Speculative Decoding.
Theorem 4.2 (Immediate Cost of the MDP).
Define $c_{1}=t_{\text{draft}}$ and $c_{2}=(t_{\text{target}}-t_{\text{draft}})$. If we set the immediate cost of the MDP to be
$c(s,{\texttt{continue}},s^{\prime})=\mathbb{I}(\exists 1\leq i\leq k+1,Y_{i}\text{ is rejected})\cdot c_{1},$
$c(s,{\texttt{stop}},s^{\prime})=\mathbb{I}(\exists 1\leq i\leq k+1,Y_{i}\text{is rejected})\cdot c_{1}+c_{2},$
then the cumulative cost of the entire trajectory equals the total inference time $T_{\text{total}}$ defined in Equation (3.1).
The theorem can be proved by invoking Lemma 3.1. For both continue and stop, we suffer a loss $c_{1}$ if the current candidate token $Y_{k+1}$ is discarded, which happens if there exists any candidate token $Y_{i}$ $(1\leq i\leq k+1)$ that is rejected. If we stop at the current step, we suffer an additional cost $c_{2}$ corresponding to the extra inference time of the target model.
Our next theorem 4.3 provides a sufficient condition for us to stop the current round of speculation and call the target model to verify the candidate tokens.
@@
the expected total cost of stop is smaller than the expected total cost of continue, where $\Delta=\Delta(\pi,x_{\text{prompt}},p,q,c_{1},c_{2})$ is a problem-specific constant.
We defer the proof of Theorem 4.3 to Appendix E.
4.3 SpecDec++ Algorithm
Motivated by Theorem 4.3, we propose SpecDec++, an adaptive speculative decoding algorithm that utilizes an additional prediction head to determine whether or not to stop the current speculation round. The additional prediction head $f_{\theta}$ is built on top of the draft model to predict the conditional probability